# Enhancing Biomedical Knowledge Discovery for Diseases: An Open-Source Framework Applied on Rett Syndrome and Alzheimer’s Disease

### v2

@@ Abstract
The ever-growing volume of biomedical publications creates a critical need for efficient knowledge discovery. In this context, we introduce an open-source end-to-end framework designed to construct knowledge around specific diseases directly from raw text. To facilitate research in disease-related knowledge discovery, we create two annotated datasets focused on Rett syndrome and Alzheimer’s disease, enabling the identification of semantic relations between biomedical entities. Extensive benchmarking explores various ways to represent relations and entity representations, offering insights into optimal modeling strategies for semantic relation detection and highlighting language models’ competence in knowledge discovery. We also conduct probing experiments using different layer representations and attention scores to explore transformers’ ability to capture semantic relations. [^1] Data and code: publicly available upon acceptance.
1 Introduction
Knowledge discovery (Wang et al., 2023; Shu and Ye, 2023) is a pivotal research domain due to the surge in publications, which makes keeping up with new findings challenging, necessitating automated knowledge extraction and processing. Of particular concern is the biomedical literature, where updates occur with ever-accelerating frequency (Fig. 1). Despite advances in healthcare, many diseases, such as Alzheimer’s disease (AD) (Scheltens et al., 2021; Trejo-Lopez et al., 2023) and multiple sclerosis (McGinley et al., 2021; Attfield et al., 2022), lack effective cures. Additionally, over 1,200 rare disorders have limited or no cures according to the National Organization for Rare Disorders. [^2] https://rarediseases.org/rare-diseases/ Discovering new scientific insights from research papers can expedite disease understanding and accelerate cure development.
@@ 1 Introduction
This paper presents an end-to-end framework for detecting medical entities in unstructured text and annotating semantic relations, enabling automated knowledge discovery for diseases. We employ a multi-stage methodology for data acquisition, annotation, and model evaluation. The process starts with gathering relevant PubMed abstracts from PubMed to form the corpus. Entities are identified and extracted, followed by the co-occurrence graph generation that models the intra-sentence co-occurrence of the entities across the corpus. Leveraging the processed text and co-occurrence graph, an algorithm samples sentences to create gold-standard datasets. Medical experts label the semantic relations between entities within these sentences via an annotation portal. The framework’s versatility allows application across various diseases and enables expansion to encompass knowledge about symptoms, genes, and more. This study focuses on two diseases of particular research interest: Rett syndrome (RS) (Petriti et al., 2023) and AD. These diseases are selected due to their significant impact and the absence of a cure, highlighting the urgency for advancements in understanding and treatment. We introduce two curated datasets tailored for detecting semantic relations between entities in biomedical text related to RS and AD. The datasets are used for benchmarking, testing techniques for representing relations and entities and assessing language models’ capabilities in knowledge discovery. This work probes the layer outputs of transformer models (Vaswani et al., 2017) and their attention patterns to reveal their ability to implicitly capture semantic relations in biomedical text.
RS (Sandweiss et al., 2020) poses challenges due to its sporadic nature and rare expression across diverse racial groups. The disorder’s elusive nature undermines its comprehension and stresses the pressing need for a cure. Rare diseases collectively affect a substantial portion of the population, with over 30 million affected people in Europe alone (Pakter, 2024). AD is characterized by its prevalence among older populations, with millions of patients worldwide as it is the most common type of dementia (60-70% cases) (Alzheimer’s-Association, 2024). With life expectancy on the rise, the projected increase in Alzheimer’s cases accentuates the urgency of finding a cure.
@@
  • Two annotated datasets for RS and AD provide gold labels for semantic relations, aiding disease knowledge discovery research. [^3] The description of distantly supervised datasets for weakly supervised scenarios is included in Appendix H.
  • Benchmarking on the datasets examines methods for relation and entity representation, offering insights into optimal approaches for semantic relation detection and emphasizing language models’ knowledge discovery capabilities.
@@
2 Data Pipeline
@@
We focus on developing a robust data pipeline (Fig. 2) to annotate sentences with entities associated with the Unified Medical Language System (UMLS) (Bodenreider, 2004; Elkin and Brown, 2023). The first step involves the retrieval of the textual abstracts, followed by the mention extraction that includes entity detection and linking to UMLS. We construct a co-occurrence graph to highlight interconnections between entities in the text. The processed text and co-occurrence graph are then used to develop two curated datasets with precise entity annotations and semantic relations between detected entity pairs. [^4] Additional information regarding the data pipeline is incorporated in Appendix A.
@@
Figure 2: The pipeline starts with abstract retrieval using a natural language query. Next, entities are detected and linked to UMLS, followed by the co-occurrence graph generation. The final step is the dataset creation using the processed text and co-occurrence graph.
@@
Abstract retrieval. We retrieve PubMed [^5] https://pubmed.ncbi.nlm.nih.gov/ articles ids based on a query (e.g., Rett syndrome) and extract their open-access abstracts. To accomplish this, we leverage the official Entrez Programming Utilities (Kans, 2024) and the Biopython API (Cock et al., 2009) (BSD 3-Clause License), ensuring access to the vast repository of biomedical literature. After obtaining the PubMed IDs (PMIDs), we retrieve the abstracts from the specified articles and tokenize the text into sentences using NLTK (Bird et al., 2009) (Apache License 2.0).
@@
Mention extraction. MetaMapLite (Aronson, 2001; Demner-Fushman et al., 2017) (open-source BSD License) is provided by the National Library of Medicine (NLM) for extracting biomedical entities and mapping them to Concept Unique Identifiers (CUIs) within UMLS. The tool is updated every two years to incorporate the latest medical terminology and to ensure its accuracy in extraction and mapping. MetaMapLite simultaneously extracts mentions and links them to UMLS in one step, efficiently associating mentions with their corresponding CUIs. We detect a diverse range of entities, spanning 82 unique semantic types and covering a broad spectrum of biomedical concepts, including diseases, biologically active substances, anatomical structures, genes, and more. Detailed entity detection often leads to overlapping or successive entities in the text. To address this, our pipeline incorporates a merging strategy that consolidates overlapping or subsequent entities into cohesive units. For example, in the sentence: "To test norepinephrine augmentation as a potential disease-modifying therapy, we performed a biomarker-driven phase II trial of atomoxetine, a clinically-approved norepinephrine transporter inhibitor, in subjects with mild cognitive impairment due to AD.", the subsequent relevant mentions norepinephrine transporter and inhibitor are merged to one entity.
@@
Co-occurrence graph generation. We model the intra-sentence co-occurrence between the entities. Each node in the graph corresponds to a unique CUI and contains metadata including the semantic type and the list of sentence IDs where the corresponding entity is detected. An edge between two nodes signifies that the corresponding entities co-occur within the same sentence. The edge weight represents the number of times two entities co-occur in a sentence throughout the text corpus.
2.1 Dataset Creation
@@
Leveraging the extracted co-occurrence graph, we define two distinct probability distributions to select sentences for manual annotation. The first distribution $\mathcal{P}$ focuses on common pairs of co-occurred entities, with higher frequency in the co-occurrence graph resulting in a higher likelihood of sampling. The second distribution $\mathcal{IP}$ prioritizes novel/rare pairs of co-occurred entities, selecting sentences where the entities have a lower frequency in the co-occurrence graph. We sample 50% of sentences using $\mathcal{P}$ and 50% using $\mathcal{IP}$ to ensure a balance of common and potentially novel pairs of co-occurring entities in the datasets. [^6] Sentence sampling algorithm details in Appendix A.
@@
Then, we develop an annotation portal using the streamlit [^7] https://streamlit.io/ library, providing a user-friendly interface for annotators. Annotators are presented with a sentence containing two highlighted entities and are prompted to categorize the semantic relation between them. Options include positive (direct semantic connection), negative (negative semantic connection where negative words like "no" and "absence" are present), complex(semantic connection with complex reasoning), and no relation. The annotation portal offers additional functions such as sentence removal (for non-informative sentences), entity removal (for incorrect entity types or spans), and context addition (for providing additional text to aid in relation type determination). We enlist the expertise of three medical experts to ensure the accuracy and reliability of the annotation process.
@@
Table 1: Datasets: Statistics of the RS and AD datasets and their label distribution.
The result of the expert annotation yields two curated datasets. The Relation Detection dataset for Rett Syndrome (ReDReS) contains 601 sentences with 5,259 instances and 1,148 unique CUIs (Tab. 1). The inter-annotator agreement is measured using the Fleiss kappa score (McHugh, 2012), resulting in 0.6143 in the multi-class setup (4 classes) and indicating substantial agreement among annotators (Landis and Koch, 1977). In the binary setup (relation or no relation), the Fleiss kappa score is 0.7139. The Relation Detection dataset for Alzheimer’s Disease (ReDAD) comprises 641 sentences with 8,565 instances and 1,480 unique CUIs (Tab. 1). The Fleiss kappa score is 0.6403 in the multi-class setup and 0.7064 in the binary setup, showing substantial consensus among annotators. The final labels are determined through majority voting, leveraging the labels provided by each expert. Whereas the label distribution across classes is relatively balanced, the negative class is under-represented with 97 and 125 instances in ReDReS and ReDAD respectively (Tab. 1). Each dataset is randomly split into train, development, and test sets.
3 Models
Figure 3: Model Architecture of LaMReDA, LaMReDM (left), and LaMEL (right): Each model encodes the input sequence using BiomedBERT (large or base). For LaMReDA and LaMReDM, different tokens define the relation representation (A-P), passed through a linear projection layer, a dropout layer, and then a classification layer for prediction. The symbol # denotes element-wise addition and multiplication for LaMReDA and LaMReDM, respectively. For LaMEL, different tokens construct the entity representation (A-H), which are sent through a dropout layer and a linear layer to extract the projected entity representations.
@@ 3 Models
3.1 LaMEL model
LaMEL learns an embedding space optimized for relation detection (Fig. 3). As the backbone language model (LM), we opt for BiomedBERT (Gu et al., 2021; Tinn et al., 2023) (MIT License), available in both uncased base and uncased large versions. [^8] HuggingFace’s Transformers library (Wolf et al., 2020) BiomedBERT is pretrained on the PubMed corpus, making it well-suited for our task as the curated datasets consist of sentences of abstracts from PubMed papers. Leveraging BiomedBERT ensures that the model can capture the language patterns prevalent in biomedical text. Following the LM encoding, we construct the representation of each entity by extracting its contextualized embedding $E_{i}$ corresponding to each entity $e_{i}$ from the encoded sequence. Subsequently, the entity representations are projected to the embedding space using a linear layer without changing the embedding dimension. The final prediction is based on cosine similarity between the two projected entity representations. If the cosine similarity exceeds a predefined threshold, the model predicts that there is a semantic relation between the two entities. We experiment with diverse strategies for learning entity representations (Fig. 3), aiming to optimize the effectiveness of the embedding space for the relation detection task. The explored types of entity representation $E$ are:

### v3

@@ Abstract
The rapidly increasing number of biomedical publications presents a significant challenge for efficient knowledge discovery. To address this, we introduce an open-source, end-to-end framework designed to automatically extract and construct knowledge about specific diseases directly from unstructured text. To facilitate research in disease-related knowledge discovery, we create two annotated datasets focused on Rett syndrome (RS) and Alzheimer’s disease (AD), enabling the identification of semantic relations between various biomedical entities. We perform extensive benchmarking to evaluate different approaches for representing relations and entities, providing insights into optimal modeling strategies for semantic relation detection and highlighting language models’ competence in knowledge discovery. To gain a deeper understanding of the internal mechanisms of transformer models, we also conduct probing experiments, analyzing different layer representations and attention scores, to explore transformers’ ability to capture semantic relations within the text. Both the code and the datasets will be publicly available [^1] https://github.com/christos42/Enhancing-Biomedical-Knowledge-Discovery-for-Diseases, encouraging further research in biomedical knowledge discovery.
Index Terms:
Benchmarking, corpus creation, evaluation, knowledge discovery, language model probing, language resources, NLP datasets, relation detection.
[…]
=-21pt
I Introduction
Knowledge discovery [1, 2] is a pivotal research domain due to the surge in publications, which makes keeping up with new findings challenging, necessitating automated knowledge extraction and processing. Of particular concern is biomedical literature, in which updates occur with increasing frequency (Fig. 1). Despite advances in healthcare, many diseases, such as Alzheimer’s disease (AD) [3, 4], multiple sclerosis [5, 6] and muscular dystrophy [7, 8], lack effective cures. Additionally, over 1,200 rare disorders have limited or no cures according to the National Organization for Rare Disorders. [^2] https://rarediseases.org/rare-diseases/ Discovering new scientific insights from research papers can expedite the understanding of disease and accelerate cure development.
@@ I Introduction
This paper presents an end-to-end framework for detecting medical entities in unstructured text and annotating semantic relations, thereby enabling automated knowledge discovery for diseases. We employ a multi-stage methodology for the data acquisition, annotation, and model evaluation. The process starts with gathering relevant PubMed abstracts from PubMed to form the corpus. Entities are identified and extracted, followed by the co-occurrence graph generation that models the intra-sentence co-occurrence of the entities across the corpus. Leveraging the processed text and co-occurrence graph, an algorithm samples sentences to create gold-standard datasets. Medical experts label the semantic relations between entities within these sentences via an annotation portal. The framework’s versatility allows application across various diseases and enables expansion to encompass knowledge about symptoms, genes, and more. This study focuses on two diseases of particular research interest: Rett syndrome (RS) [9] and AD. These diseases are selected due to their significant impact and the absence of a cure, highlighting the urgency for advancements in understanding and treatment. We introduce two curated datasets tailored to detect semantic relations between entities in biomedical text related to RS and AD. The datasets are used for benchmarking, testing techniques for representing relations and entities, and assessing language models’ capabilities in knowledge discovery. This work probes the layer outputs of transformer models [10] and their attention patterns to reveal their ability to implicitly capture semantic relations in biomedical text.
RS [11] poses challenges due to its sporadic nature and rare expression across diverse racial groups. The elusive nature of this disorder undermines its comprehension and stresses the pressing need for a cure. Rare diseases collectively affect a substantial portion of the population, with over 30 million people affected in Europe alone [12]. AD is characterized by its prevalence among older populations, with millions of patients worldwide as it is the most common type of dementia (60-70% of cases) [13]. With life expectancy on the rise, the projected increase in Alzheimer’s disease’s cases accentuates the urgency of finding a cure.
@@
  • Two annotated datasets for RS and AD provide gold labels for semantic relations, aiding disease knowledge discovery research.
  • Benchmarking on the datasets examines methods for relation and entity representation, offering insights into optimal approaches for semantic relation detection and emphasizing the knowledge discovery capabilities of language models.
@@
II Related Work
Information Extraction Datasets. Several biomedical datasets aim to enhance Information Extraction (IE) system development [14, 15, 16, 17], typically focusing on one or a few entity types and their interactions. AIMed [18], BioInfer [19], and BioCreative II PPI IPS [20] formulate protein-protein interactions. The chemical-protein and chemical-disease interactions are modeled by DrugProt [21] and BC5CDR [22], respectively. ADE [23], DDI13 [24], and n2c2 2018 ADE [25] include drug-ADE (adverse drug effect) and drug-drug interactions. EMU [26], GAD [27] and RENET2 [28] contain relations between genes and diseases. N-ary [29] incorporates drug-gene mutation interactions. The event extraction task is illustrated by GE09 [30], GE11 [31], and CG [32]. DDAE [33] includes disease-disease associations. BioRED [34] focuses on document-level relations for various entities. Unlike these datasets, the datasets of our paper focus on RS and AD, including entities of up to 82 different semantic types, and model the semantic relation between them.
Knowledge Discovery. [35] presents PREDICT, a method for ranking potential drug-disease associations to predict drug indications. [36] releases AlzKB, a heterogeneous graph knowledge base for AD, constructed using external data sources and describing various medical entities (e.g., chemicals and genes). Other graph-based efforts model knowledge around AD for tasks such as drug repurposing [37, 38, 39], gene identification [40], or as general knowledge repositories [41]. Recently, the study by [42] focuses on a recommendation system that suggests medical articles based on diagnoses in discharge summaries, which involves semantic similarity modeling between summaries and scientific publications. [43] introduces MLTE, a feature fusion approach for semantic textual similarity analysis, which aims to measure the similarity between pairs of clinical texts.
Another paradigm for knowledge discovery is the open information extraction (OIE) setup [44, 45], which faces challenges such as data consistency, performance evaluation, and semantic drift [46]. Research efforts [47, 48, 49, 50, 51] aim to address these problems and extract knowledge with little or no supervision. Advances in literature-based discovery [52, 53] try to identify novel medical entity relations using graph-based [54, 55], machine learning [56, 57], and co-occurrence methods [58, 59]. [60] stresses the potential of large language models (LLMs) to summarize, simplify, and synthesize medical evidence [61, 62, 63], suggesting that LLMs may have encoded biomedical knowledge [64]. To exploit this potential, we explore the construction of LM representations for knowledge discovery. To the best of our knowledge, no systematic approach assembles knowledge about RS. Unlike previous work, we introduce a, in principle, disease-agnostic framework, to acquire knowledge about RS and AD starting from raw text.
III Data Pipeline
@@
We focus on developing a robust data pipeline (Fig. 2) to annotate sentences with entities associated with the Unified Medical Language System (UMLS) [65, 66]. The first step involves the retrieval of textual abstracts, followed by mention extraction, which includes entity detection and linking to UMLS. We construct a co-occurrence graph to highlight the interconnections between the entities in the text. The processed text and co-occurrence graph are then used to develop two curated datasets with precise entity annotations and semantic relations between the detected entity pairs. Appendix A provides additional information on the data pipeline. We state that the selection of the open-source tools is based on their wide adoption by the research community, their accessibility, and their suitability for biomedical text processing.
@@
Figure 2: The pipeline starts with abstract retrieval using a natural language query. Next, entities are detected and linked to UMLS, followed by the co-occurrence graph generation. The final step is the dataset creation using the processed text (abstract retrieval and mention extraction steps) and co-occurrence graph.
@@
Abstract retrieval. We retrieve the PubMed [^3] https://pubmed.ncbi.nlm.nih.gov/ articles ids based on a query (e.g., Rett syndrome) and extract their open-access abstracts. To accomplish this, we leverage the official Entrez Programming Utilities [67] and the Biopython API [68] (BSD 3-Clause License), ensuring access to a vast repository of biomedical literature. After obtaining the PubMed IDs (PMIDs), we retrieve abstracts from the specified articles and tokenize the text into sentences using NLTK [69] (Apache License 2.0). To facilitate effective abstract retrieval, we implement an iterative approach to circumvent the API’s limitation of retrieving only 10,000 article IDs per query. This iterative process enables us to access a comprehensive set of PubMed IDs (PMIDs) related to the query.
@@
Mention extraction. MetaMapLite [70, 71] (open-source BSD License) is provided by the National Library of Medicine (NLM) for extracting biomedical entities and mapping them to Concept Unique Identifiers (CUIs) within UMLS. The tool is updated every two years to incorporate the latest medical terminology and to ensure its accuracy in extraction and mapping. MetaMapLite simultaneously extracts mentions and links them to UMLS in one step, efficiently associating the mentions with their corresponding CUIs. We detect a diverse range of entities, spanning 82 unique semantic types and covering a broad spectrum of biomedical concepts, including diseases, biologically active substances, anatomical structures, genes, and more (Table I). Detailed entity detection often leads to overlapping or successive entities in text. To address this issue, our pipeline incorporates a heuristic merging strategy that consolidates overlapping or subsequent entities into cohesive units. For example, in the sentence: ”To test norepinephrine augmentation as a potential disease-modifying therapy, we performed a biomarker-driven phase II trial of atomoxetine, a clinically-approved norepinephrine transporter inhibitor, in subjects with mild cognitive impairment due to AD.”, the subsequent relevant mentions norepinephrine transporter and inhibitor are merged into one entity.
TABLE I: List of the 82 semantic types of the MetaMapLite-based pipeline.
Semantic Types
Bacterium Body Substance Body System Behavior Biologic Function
Plant Organism Population Group Receptor Reptile
Virus Vitamin Vertebrate
@@
Co-occurrence graph generation. We model the intra-sentence co-occurrence between the entities (an example of a co-occurrence subgraph is presented in Fig. 3). Each node in the graph corresponds to a unique CUI and contains metadata including the semantic type and list of sentence IDs where the corresponding entity is detected. An edge between two nodes signifies that corresponding entities co-occur within the same sentence. The edge weight represents the number of times two entities co-occur in a sentence throughout the text corpus.
Figure 3: Visualization of a subgraph of the co-occurrence graph for the Rett Syndrome corpus. Each node corresponds to a unique CUI with the related textual description and contains the semantic type. The edge label includes the number of times two entities co-occur in a sentence and a list with the sentence IDs where the connected entities are detected.
@@
Dataset creation. Leveraging the extracted co-occurrence graph, we define two distinct probability distributions to select sentences for manual annotation. The first distribution $\mathcal{P}$ focuses on common pairs of co-occurred entities, with higher frequency in the co-occurrence graph resulting in a higher likelihood of sampling. The second distribution $\mathcal{IP}$ prioritizes novel/rare pairs of co-occurred entities, selecting sentences where the entities have a lower frequency in the co-occurrence graph. We sample 50% of sentences using $\mathcal{P}$ and 50% using $\mathcal{IP}$ to ensure a balance of common and potentially novel pairs of co-occurring entities in the datasets. The sentence sampling algorithm is detailed in Appendix A.
@@
Then, we develop an annotation portal using the streamlit [^4] https://streamlit.io/ library, providing a user-friendly interface for annotators. Annotators are presented with a sentence containing two highlighted entities and are prompted to categorize the semantic relation between them. Options include positive (direct semantic connection), negative (negative semantic connection where negative words like ”no” and ”absence” are present), complex, and no relation (absence of semantic relation). In collaboration with clinical experts, we identify instances where entities exhibit a semantic connection without a clear positive or negative semantic relation. To capture these complexities, we introduce the ”complex” relation type, enhancing the potential for in-depth analysis. The annotation portal offers additional functions such as sentence removal (for non-informative sentences), entity removal (for incorrect entity types or spans), and context addition (for providing additional text to aid in relation type determination). We enlist the expertise of three medical experts to ensure the accuracy and reliability of the annotation process.
TABLE II: Datasets: Statistics of ReDReS and ReDAD and their label distribution.
@@
Figure 4: Graphical example of the ReDAD and ReDReS datasets. Each node corresponds to an entity with a textual description and semantic type. The edge label includes the annotated relation type and the sentence ID where the connected entities are detected.
The result of the expert annotation yields two curated datasets. The Relation Detection dataset for Rett Syndrome (ReDReS) contains 601 sentences with 5,259 instances and 1,148 unique CUIs (Table II). The inter-annotator agreement is measured using the Fleiss kappa score [72], resulting in 0.6143 in the multi-class setup (4 classes) and indicating substantial agreement among annotators [73]. In the binary setup (relation or no relation), where the classes positive, complex, negative are grouped under the relation class, the Fleiss kappa score is 0.7139. The Relation Detection dataset for Alzheimer’s Disease (ReDAD) comprises 641 sentences with 8,565 instances and 1,480 unique CUIs (Table II). The Fleiss kappa score is 0.6403 in the multi-class setup and 0.7064 in the binary setup, showing substantial consensus among annotators. The final labels are determined through majority voting, leveraging the labels provided by each expert. Whereas the label distribution across classes is relatively balanced, the negative class is under-represented with 97 and 125 instances in ReDReS and ReDAD respectively (Table II). Each dataset is randomly split into train, development, and test sets. A graphical example of the datasets is presented in Fig. 4.
IV Models
@@
Figure 5: Model Architecture of LaMReDA, LaMReDM (left), and LaMEL (right): Each model encodes the input sequence using BiomedBERT (large or base). For LaMReDA and LaMReDM, different tokens define the relation representation (A-P), passed through a linear projection layer, a dropout layer, and then a classification layer for prediction. The symbol # denotes element-wise addition and multiplication for LaMReDA and LaMReDM, respectively. For LaMEL, different tokens construct the entity representation (A-H), which are sent through a dropout layer and a linear layer to extract the projected entity representations. The symbols ; and * define the concatenation and the element-wise multiplication, respectively, for the LaMEL model.
IV-A LaMEL model
LaMEL (Fig. 5) learns an embedding space optimized for relation detection (Fig. 5). As the backbone language model (LM), we opt for BiomedBERT [74, 75] (MIT License), which is available in both uncased base and uncased large versions, accessed via HuggingFace’s Transformers library [76]. BiomedBERT is pre-trained on the PubMed corpus, making it well-suited for our task as the curated datasets consist of sentences of abstracts from PubMed papers. Leveraging BiomedBERT ensures that the model can capture language patterns prevalent in biomedical texts. Following LM encoding, we construct the representation of each entity by extracting its contextualized embedding $E_{i}$ corresponding to each entity $e_{i}$ from the encoded sequence. Subsequently, entity representations are projected onto the embedding space using a linear layer without changing the embedding dimension. The final prediction is based on the cosine similarity between the two projected entity representations. If the cosine similarity exceeds a predefined threshold, the model predicts a semantic relation between the two entities. This approach leverages the geometric properties of the embedding space to determine the semantic similarity between entities. We experiment with diverse strategies for learning entity representations (Fig. 5), aiming to optimize the effectiveness of the embedding space for the relation detection task. The explored types of entity representation $E$ are: